Dave was on site all weekend working on cleaning and painting one of the bodyside radiator grilles.
Ian and Pete with some assistance from Dave continued working on the ex-50008 engine replacing some of the head covers.
Steve and Sarah assessed traction motor contactor arc-chutes so spares and paxolin to replace missing parts can be ordered.
The long process of verifying all the cables from the electrical cubicle to equipment in 50029 was started. Each cable was to be tested with a multi-meter to confirm continuity. The first cables checked were those from terminal bar TB6 with blower No1 cables testing out ok. Problems were then encountered as we checked Blower No2 but could get no reading and eventually realised this was because the cables and flexible conduit had been cut about an inch from the connection box.
As it is too difficult to pull any slack through we have decided to dismantle the conduits with a view to reinstalling via a shorter route up the front of the blower negating the need to replace the cables.